OpenStudy (anonymous):

Simplify (x^1/3)^3/5

OpenStudy (anonymous):

I think it's x^1/2 but I'm not sure.

Parth (parthkohli):

\( \color{Black}{\Rightarrow (a^x)^y = a^{x\cdot y} }\)

OpenStudy (anonymous):

multiply the exponents...

OpenStudy (saifoo.khan):

\[\Large x^{ \frac 13 \times \frac 35}\]

OpenStudy (anonymous):

round and round.... medals...

Parth (parthkohli):

lol exactly

OpenStudy (anonymous):

So it's x^1/5! Right?

OpenStudy (saifoo.khan):

Right.

Parth (parthkohli):

\( \color{Black}{\Rightarrow {1 \over 3} \times {3 \over 5} }\) This will be the exponent. Simplify.

Parth (parthkohli):

Yes, CORRECTO!

OpenStudy (anonymous):

Okay! Thank you. :) I just had the wrong idea.

OpenStudy (anonymous):

round and round .... answers...

Parth (parthkohli):

lolol

OpenStudy (anonymous):

good work team!

Parth (parthkohli):

99 team B)

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

round and round.... comments.... round and round... dpanic's frog's eyes.....

OpenStudy (anonymous):

cooler ---> B)

Parth (parthkohli):

dpanic? lol

OpenStudy (saifoo.khan):

\[\begin{array}l\color{red}{\normalsize\text{L}}\color{orange}{\normalsize\text{e}}\color{#9c9a2e}{\normalsize\text{t}}\color{green}{\normalsize\text{'}}\color{blue}{\normalsize\text{s}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{a}}\color{red}{\normalsize\text{l}}\color{orange}{\normalsize\text{l}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{b}}\color{blue}{\normalsize\text{e}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{c}}\color{red}{\normalsize\text{o}}\color{orange}{\normalsize\text{o}}\color{#9c9a2e}{\normalsize\text{l}}\color{green}{\normalsize\text{.}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{B}}\color{purple}{\normalsize\text{)}}\color{red}{\normalsize\text{}}\end{array}\]

Parth (parthkohli):

You got that from the Amistre's thing ^

OpenStudy (anonymous):

sunglasses emoticon... B) i'm cool man...

OpenStudy (saifoo.khan):

No, you aren't a man. lol. you are cool woman.

OpenStudy (anonymous):

dooh!

OpenStudy (saifoo.khan):

Haha.

Parth (parthkohli):

http://i.imgur.com/MSBcG.jpg

OpenStudy (anonymous):

yeah parth!

Parth (parthkohli):

:P